# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ Appendix B Validity of Remark 1 within our linear coding framework
Let (${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) be the solution tuple of the optimization in (25). Let’s consider $\tilde{P}>P$. We can always find $\tilde{\bf g}_{1}=(1+\epsilon){\bf g}_{1}$, where $\epsilon>0$ can be chosen to satisfy both $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq N\tilde{P}$ and $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq N\tilde{P}$ in (16)-(17). Then, the solution tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) yields larger $\text{SNR}_{1}$ in (23), leading to smaller $\mathcal{E}_{1}(\text{SNR}_{1})$. However, $\mathcal{E}_{2}(\text{SNR}_{2})$ remains the same here since $\text{SNR}_{2}$ is not a function of $\tilde{\bf g}_{1}$. Thus, $\mathcal{E}_{1}(\text{SNR}_{1})+\mathcal{E}_{2}(\text{SNR}_{2})$ decreases.
Appendix C Proof of Lemma 1
@@
We show this by contradiction. Suppose that $h(\delta,\eta_{2})$ is not an increasing function of $\delta$. This implies that there exist $\delta_{1}$ and $\delta_{2}$, such that $\delta_{1}<\delta_{2}$ and $h(\delta_{1},\eta_{2})\geq h(\delta_{2},\eta_{2})$. Let us denote the optimal solution tuple that yields $h(\delta_{2},\eta_{2})$ and satisfies the constraints, $\text{SNR}_{1}=\delta_{2}$ and $\text{SNR}_{2}=\eta_{2}$, as (${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$). Assuming that ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$ are fixed, we can find $\tilde{\bf g}_{1}=(1-\epsilon){\bf g}_{1}$, where $\epsilon>0$ is chosen to satisfy $\text{SNR}_{1}=\delta_{1}<\delta_{2}$ in (23). The tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) satisfies $\text{SNR}_{2}=\eta_{2}$ since $\text{SNR}_{2}$ is not a function of ${\bf g}_{1}$. Furthermore, this tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) yields smaller $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ in (16) and $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ in (17), since $\|\tilde{\bf g}_{1}\|^{2}<\|{\bf g}_{1}\|^{2}$ and $\|{\bf F}_{2}\tilde{\bf g}_{1}\|^{2}<\|{\bf F}_{2}{\bf g}_{1}\|^{2}$. That is, the tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) yields a smaller objective value of ${\max}\big{\{}\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$, say $\nu$, than the one obtained by (${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$), i.e., $\nu<h(\delta_{2},\eta_{2})$.
@@
Appendix D Proof of Proposition 1
@@
Proof of Statement (ii): First, (${\bf g}_{1}$, ${\bf F}_{1}$, $\bar{\bf g}_{2}$, $\bar{\bf F}_{2}$) yields a smaller or an equal transmit power of $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ since
$\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ $=\|({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf g}_{1}\|^{2}+\|{\bf F}_{2}({\bf I}+{\bf F}_{1}{\bf F}_{2})\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}$
@@
Note that $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ in (16) are not dependent on $f_{2,N}$ since ${\bf F}_{1}{\bf F}_{2}$ does not include $f_{2,N}$.
Therefore, when $f_{2,N}=0$, we can always obtain a smaller or an equal objective value of $\alpha\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ in (29), while satisfying the SNR constraints. It is thus optimal to set $f_{2,N}=0$ in (29) with ${\bf F}_{2}\in\mathcal{F}_{2}$. ∎
Appendix E Proof of Conjecture 1 When $N=3$
@@
We note that $\min\{\alpha\sigma_{1}^{2},(1-\alpha)\sigma_{2}^{2}\}\leq(1-\alpha)\sigma_{2}^{2}$ for any $\alpha\in(0,1)$. In the special case with $N=3$, we will show that $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$ for any ${\bf F}_{1}$ and ${\bf F}_{2}\in\mathcal{F}_{2}$ (in the form of (30)). We first rewrite ${\bf B}=(1-\alpha)\sigma_{2}^{2}{\bf I}+{\bf C}$ where ${\bf C}=(1-\alpha)\sigma_{1}^{2}{\bf F}_{2}{\bf F}_{2}^{\top}+{\bf Q}_{2}^{1/2}\big{(}\alpha{\bf F}_{1}^{\top}{\bf F}_{1}+(1-\alpha)({\bf F}_{2}{\bf F}_{1}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf F}_{1})\big{)}{\bf Q}_{2}^{1/2}$. Then, showing $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$ is equivalent to showing $\nu_{\min}[{\bf C}]=0$. First, since ${\bf F}_{2}{\bf F}_{1}={\bf 0}$ due to $f_{2,3}=0$ from Proposition 1, we have
${\bf C}=(1-\alpha)\sigma_{1}^{2}{\bf F}_{2}{\bf F}_{2}^{\top}+\alpha{\bf Q}_{2}^{1/2}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2}^{1/2}.$ (83)
In (83), ${\bf F}_{2}{\bf F}_{2}^{\top}$ and ${\bf Q}_{2}^{1/2}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2}^{1/2}$ are positive semidefinite (PSD), respectively, since each of them is a form of a matrix multiplied with its own transpose. Since $(1-\alpha)\sigma_{1}^{2}>0$ and $\alpha>0$, $(1-\alpha)\sigma_{1}^{2}{\bf F}_{2}{\bf F}_{2}^{\top}$ and $\alpha{\bf Q}_{2}^{1/2}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2}^{1/2}$ in (83) are PSD, respectively. Thus, the summation of the two PSD matrices is also PSD, i.e., ${\bf C}$ is PSD. This implies that $\nu_{\min}[{\bf C}]\geq 0$. The remaining part to claim $\nu_{\min}[{\bf C}]=0$ is to show that ${\bf C}$ has an eigenvalue of 0. Since the last ($N$-th) column of ${\bf C}$ is a zero vector, ${\bf C}$ has an eigenvalue of 0 with the corresponding eigenvector $[0,0,1]^{\top}$, i.e., ${\bf C}[0,0,1]^{\top}=0\cdot[0,0,1]^{\top}$. We then have $\nu_{\min}[{\bf C}]=0$, which leads to $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$. ∎
Appendix F Proof of Proposition 2
@@
We have a lower bound of the objective function in (32) as ${\bf q}_{2}^{\top}{\bf B}{\bf q}_{2}\geq\nu_{\min}[{\bf B}]\|{\bf q}_{2}\|^{2}.$ From Conjecture 1, we have $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$ when $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$. Then, we have the lower bound as ${\bf q}_{2}^{\top}{\bf B}{\bf q}_{2}\geq(1-\alpha)\sigma_{2}^{2}\|{\bf q}_{2}\|^{2}.$ Here, ${\bf q}_{2}^{\star}=[0,...,0,\sqrt{\eta_{2}}]^{\top}$ satisfies the lower bound with $\|{\bf q}^{\star}_{2}\|^{2}=\eta_{2}$, which can be easily shown by the fact that all the entries in the last column and row of ${\bf B}$ are zeros except the last diagonal entry is $(1-\alpha)\sigma_{2}^{2}$ due to $f_{2,N}=0$ from Proposition 1. In other words, ${\bf q}_{2}^{\star}$ is an optimal solution of (32). We then have the optimal solution for (31) as ${\bf g}_{2}^{\star}={\bf Q}_{2}^{1/2}{\bf q}_{2}^{\star}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$, since $f_{2,N}=0$ from Proposition 1. ∎
Appendix G Proof of Lemma 2

### v2

@@ Appendix D Validity of Remark 2 within our linear coding framework
Let (${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) be the solution tuple of the optimization in (41). Let’s consider $\tilde{P}>P$. We can always find $\tilde{\bf g}_{1}=(1+\epsilon){\bf g}_{1}$, where $\epsilon>0$ can be chosen to satisfy both $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq N\tilde{P}$ and $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq N\tilde{P}$ in (16)-(17). Then, the solution tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) yields larger $\text{SNR}_{1}$ in (39), leading to smaller $\mathcal{E}_{1}(\text{SNR}_{1})$. However, $\mathcal{E}_{2}(\text{SNR}_{2})$ remains the same here since $\text{SNR}_{2}$ is not a function of $\tilde{\bf g}_{1}$. Thus, $\mathcal{E}_{1}(\text{SNR}_{1})+\mathcal{E}_{2}(\text{SNR}_{2})$ decreases.
Appendix E Proof of Lemma 1
Proof.
Denote a tuple of variables as $s=({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$. To indicate the dependency on $s$, we denote the SNR values as $\text{SNR}_{1}(s)$ and $\text{SNR}_{2}(s)$, and on the power value as $P_{i}(s)=\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{i}\|^{2}\big{]}$, for $i=1,2$. We first define the set of optimal solutions of (41) as $\mathcal{S}^{\star}$. For any optimal solution $s^{\star}\in\mathcal{S}^{\star}$ for (41), the minimum objective value is $\mathcal{E}_{1}(\text{SNR}_{1}(s^{\star}))+\mathcal{E}_{2}(\text{SNR}_{2}(s^{\star}))=w^{\star}$, and the power constraint is satisfied as $P_{i}(s^{\star})\leq NP$ for $i=1,2$ or equivalently as $\max\{P_{1}(s^{\star}),P_{2}(s^{\star})\}\leq NP$. We define the set of optimal solutions in (45) as $\mathcal{\bar{S}}$. For any optimal solution $\bar{s}\in\mathcal{\bar{S}}$, the minimum objective value in (45) is $\mathcal{E}_{1}(\text{SNR}_{1}(\bar{s}))+\mathcal{E}_{2}(\text{SNR}_{2}(\bar{s}))=\bar{w}$. Recall that $\mathcal{S}$ is the feasible solution set of (45), and any feasible solution $s\in\mathcal{S}$ satisfies the power constraint, $\max\{P_{1}(s),P_{2}(s)\}\leq NP$, by its definition. Therefore, any optimal solution $\bar{s}\in\mathcal{\bar{S}}\subset\mathcal{S}$ satisfies the power constraint.
Claim. We claim that $\mathcal{S}^{\star}=\mathcal{\bar{S}}$. We prove this by showing the two following statements: (i) if $s_{0}\in\mathcal{S}^{\star}$, then $s_{0}\in\mathcal{\bar{S}}$, and (ii) if $s_{0}\in\mathcal{\bar{S}}$, then $s_{0}\in\mathcal{S}^{\star}$.
Proof of statement (i) by contrapositive. The contrapositive of statement (i) is that, if $s_{0}\notin\mathcal{\bar{S}}$, then $s_{0}\notin\mathcal{S}^{\star}$. Suppose $s_{0}\notin\mathcal{\bar{S}}$. Then, we have either of the two following cases.
Case (i-a). $s_{0}$ does not satisfy the power constraint, i.e., $P_{i}(s_{0})>NP$ for some $i\in\{1,2\}$.
Case (i-b). $s_{0}$ does not minimize the sum-error value in (45) while satisfying the power constraints, i.e., $\mathcal{E}_{1}(\text{SNR}_{1}(s_{0}))+\mathcal{E}_{2}(\text{SNR}_{2}(s_{0}))=w^{o}>\bar{w}$ and $P_{i}(s_{0})\leq NP$, $i=1,2$.
First, we consider Case (i-a). Since $s_{0}$ does not satisfy the power constraint, it is not a solution of (41). Therefore, it cannot be an optimal solution for (41), i.e., $s_{0}\notin\mathcal{S}^{\star}$. Next, we consider Case (i-b). We first note that $\bar{s}\in\mathcal{\bar{S}}$ is a feasible solution of (41), since $\bar{s}$ satisfies the power constraints. It is obvious that $\bar{w}\geq w^{\star}$ since the feasible solution space of (45) is a subset of that of (41). From Case (i-b), we have $w^{o}>\bar{w}$. Combining the previous two results, we arrive at $w^{o}>w^{\star}$. This shows that $s_{0}$ does not yield the lowest sum-error value in (41), and thus it is not an optimal solution of (41), i.e., $s_{0}\notin\mathcal{S}^{\star}$. Therefore, for either of the cases, $s_{0}\notin\mathcal{S}^{\star}$.
Proof of statement (ii) by contrapositive. The contrapositive of statement (ii) is that, if $s_{0}\notin\mathcal{S}^{\star}$, then $s_{0}\notin\mathcal{\bar{S}}$. Suppose $s_{0}\notin\mathcal{S}^{\star}$. Then, we have either of the two following cases.
Case (ii-a). $s_{0}$ does not satisfy the power constraint, i.e., $P_{i}(s_{0})>NP$ for some $i\in\{1,2\}$.
Case (ii-b). $s_{0}$ does not minimize the sum-error value in (41) while satisfying the power constraints, i.e., $\mathcal{E}_{1}(\text{SNR}_{1}(s_{0}))+\mathcal{E}_{2}(\text{SNR}_{2}(s_{0}))=w^{o}>w^{\star}$ and $P_{i}(s_{0})\leq NP$, $i=1,2$.
First, we consider Case (ii-a). Since $s_{0}$ does not satisfy the power constraint, it is not a solution of (45), i.e., $s_{0}\notin\mathcal{\bar{S}}$. Next, we consider Case (ii-b). We note that $s^{\star}$ is a feasible solution of (44) when $\eta_{1}=\text{SNR}_{1}(s^{\star})$ and $\eta_{2}=\text{SNR}_{2}(s^{\star})$. Also, $s^{\star}$ satisfies the power constraints, $P_{i}(s^{\star})\leq NP$, $i=1,2$. Therefore, $s^{\star}$ is a feasible solution of (45), i.e., $s^{\star}\in\mathcal{S}$. From Case (ii-b), we have $w^{o}>w^{\star}$. This means that $s_{0}$ yields a higher objective value than an existing feasible solution $s^{\star}$ of (45), and thus $s_{0}$ cannot be an optimal solution of (45), i.e., $s_{0}\notin\mathcal{\bar{S}}$. For either of the cases, $s_{0}\notin\mathcal{\bar{S}}$. ∎
Appendix F Proof of Lemma 2
@@
We show this by contradiction. Suppose that $h(\delta,\eta_{2})$ is not an increasing function of $\delta$. This implies that there exist $\delta_{1}$ and $\delta_{2}$, such that $\delta_{1}<\delta_{2}$ and $h(\delta_{1},\eta_{2})\geq h(\delta_{2},\eta_{2})$. Let us denote the optimal solution tuple that yields $h(\delta_{2},\eta_{2})$ and satisfies the constraints, $\text{SNR}_{1}=\delta_{2}$ and $\text{SNR}_{2}=\eta_{2}$, as (${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$). With fixed ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, we can find $\tilde{\bf g}_{1}=(1-\epsilon){\bf g}_{1}$, where $10<\epsilon<$ 1 is chosen to satisfy $\text{SNR}_{1}=\delta_{1}<\delta_{2}$ in (39). The tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) satisfies $\text{SNR}_{2}=\eta_{2}$ since $\text{SNR}_{2}$ is not a function of ${\bf g}_{1}$. Furthermore, this tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) yields smaller $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ in (16) and $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ in (17), since $\|\tilde{\bf g}_{1}\|^{2}<\|{\bf g}_{1}\|^{2}$ and $\|{\bf F}_{2}\tilde{\bf g}_{1}\|^{2}<\|{\bf F}_{2}{\bf g}_{1}\|^{2}$. That is, the tuple ($\tilde{\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$) yields a smaller objective value of ${\max}\big{\{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$, say $\nu$, than the one obtained by (${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, ${\bf F}_{2}$), i.e., $\nu<h(\delta_{2},\eta_{2})$.
@@
Appendix G Proof of Proposition 1
@@
Proof of Statement (ii): First, (${\bf g}_{1}$, ${\bf F}_{1}$, $\bar{\bf g}_{2}$, $\bar{\bf F}_{2}$) yields a smaller or an equal transmit power of $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ since
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ $=\|({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf g}_{1}\|^{2}+\|{\bf F}_{2}({\bf I}+{\bf F}_{1}{\bf F}_{2})\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}$
@@
Note that $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ in (16) are not dependent on $f_{2,N}$ since ${\bf F}_{1}{\bf F}_{2}$ does not include $f_{2,N}$.
Therefore, when $f_{2,N}=0$, we can always obtain a smaller or an equal objective value of $\alpha\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}$ $\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ in (46), while satisfying the SNR constraints. It is thus optimal to set $f_{2,N}=0$ in (46) with ${\bf F}_{2}\in\mathcal{F}_{2}$. ∎
Appendix H Proof of Proposition 2
Proof.
We first prove the upper bound, $\nu_{\min}[{\bf B}]\leq(1-\alpha)\sigma_{2}^{2}$. We rewrite ${\bf B}$ in (50) as ${\bf B}=\alpha{\bf Q}_{2,\text{sqrt}}{\bf F}_{1}^{\top}{\bf F}_{1}$ ${\bf Q}_{2,\text{sqrt}}+(1-\alpha){\bf Q}_{2}+(1-\alpha){\bf Q}_{2,\text{sqrt}}\big{(}{\bf F}_{2}{\bf F}_{1}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf F}_{1}\big{)}{\bf Q}_{2,\text{sqrt}}$. Then, we obtain ${\bf B}=(1-\alpha)\sigma_{2}^{2}{\bf I}+{\bf C}$ by using ${\bf Q}_{2}={\bf F}_{2}{\bf F}_{2}^{\top}\sigma_{1}^{2}+\sigma_{2}^{2}{\bf I}$ from (21) and defining ${\bf C}=(1-\alpha)\sigma_{1}^{2}{\bf F}_{2}{\bf F}_{2}^{\top}+{\bf Q}_{2,\text{sqrt}}\big{(}\alpha{\bf F}_{1}^{\top}{\bf F}_{1}+(1-\alpha)({\bf F}_{2}{\bf F}_{1}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf F}_{1})\big{)}{\bf Q}_{2,\text{sqrt}}$. By the Weyl’s inequality, $\nu_{\min}[{\bf B}]\leq(1-\alpha)\sigma_{2}^{2}+\nu_{\min}[{\bf C}]$. Showing $\nu_{\min}[{\bf B}]\leq(1-\alpha)\sigma_{2}^{2}$ is equivalent to showing $\nu_{\min}[{\bf C}]\leq 0$. We note that all the entries at the last column and row of ${\bf C}$ are zeros for any strictly lower triangular matrix ${\bf F}_{1}$ and any ${\bf F}_{2}\in\mathcal{F}_{2}$. Since the last ($N$-th) column of ${\bf C}$ is a zero vector, ${\bf C}$ has an eigenvalue of 0 with the corresponding eigenvector $[0,...,0,1]^{\top}$, i.e., ${\bf C}[0,...,0,1]^{\top}=0\cdot[0,...,0,1]^{\top}$. Since ${\bf C}$ has an eigenvalue of 0, the smallest eigenvalue of ${\bf C}$ should be less than or equal to 0, i.e., $\nu_{\min}[{\bf C}]\leq 0$. Then, $\nu_{\min}[{\bf B}]\leq(1-\alpha)\sigma_{2}^{2}$.
Next, we prove the lower bound, $\nu_{\min}[{\bf B}]\geq 0$. We first define the terms of ${\bf B}$ in (50) as ${\bf B}_{1}=\alpha{\bf Q}_{2,\text{sqrt}}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2,\text{sqrt}}$ and ${\bf B}_{2}=(1-\alpha){\bf Q}_{2,\text{sqrt}}({\bf I}+{\bf F}_{2}{\bf F}_{1})^{\top}({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf Q}_{2,\text{sqrt}}$, so that ${\bf B}={\bf B}_{1}+{\bf B}_{2}$. Note that ${\bf B}_{1}$ and ${\bf B}_{2}$, are positive semidefinite (PSD), respectively. Therefore, their sum, ${\bf B}$, is also PSD. This implies that $\nu_{\min}[{\bf B}]\geq 0$. ∎
Appendix I Proof of Conjecture 1 When $N=3$
@@
We will show $\min\{\alpha\sigma_{1}^{2},(1-\alpha)\sigma_{2}^{2}\}\leq\nu_{\min}[{\bf B}]$. By understanding that $\min\{\alpha\sigma_{1}^{2},(1-\alpha)\sigma_{2}^{2}\}\leq(1-\alpha)\sigma_{2}^{2}$, in the special case with $N=3$, we will instead show that $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$ for any strictly lower triangular ${\bf F}_{1}$ and any ${\bf F}_{2}\in\mathcal{F}_{2}$. We first rewrite ${\bf B}=(1-\alpha)\sigma_{2}^{2}{\bf I}+{\bf C}$ where ${\bf C}=(1-\alpha)\sigma_{1}^{2}{\bf F}_{2}{\bf F}_{2}^{\top}+{\bf Q}_{2,\text{sqrt}}\big{(}\alpha{\bf F}_{1}^{\top}{\bf F}_{1}+(1-\alpha)({\bf F}_{2}{\bf F}_{1}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}+{\bf F}_{1}^{\top}{\bf F}_{2}^{\top}{\bf F}_{2}{\bf F}_{1})\big{)}{\bf Q}_{2,\text{sqrt}}$. Then, showing $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$ is equivalent to showing $\nu_{\min}[{\bf C}]=0$. First, since ${\bf F}_{2}{\bf F}_{1}={\bf 0}$ due to $f_{2,3}=0$ from Proposition 1, we have
${\bf C}=(1-\alpha)\sigma_{1}^{2}{\bf F}_{2}{\bf F}_{2}^{\top}+\alpha{\bf Q}_{2,\text{sqrt}}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2,\text{sqrt}}.$ (138)
In (138), ${\bf F}_{2}{\bf F}_{2}^{\top}$ and ${\bf Q}_{2,\text{sqrt}}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2,\text{sqrt}}$ are positive semidefinite (PSD), respectively, since each of them is a form of a matrix multiplied with its own transpose. Since $(1-\alpha)\sigma_{1}^{2}>0$ and $\alpha>0$, $(1-\alpha)\sigma_{1}^{2}{\bf F}_{2}{\bf F}_{2}^{\top}$ and $\alpha{\bf Q}_{2,\text{sqrt}}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2,\text{sqrt}}$ in (138) are PSD, respectively. Thus, the summation of the two PSD matrices is also PSD, i.e., ${\bf C}$ is PSD. This implies that $\nu_{\min}[{\bf C}]\geq 0$. The remaining part to claim $\nu_{\min}[{\bf C}]=0$ is to show that ${\bf C}$ has an eigenvalue of 0. Since the last ($N$-th) column of ${\bf C}$ is a zero vector, ${\bf C}$ has an eigenvalue of 0 with the corresponding eigenvector $[0,0,1]^{\top}$, i.e., ${\bf C}[0,0,1]^{\top}=0\cdot[0,0,1]^{\top}$. We then have $\nu_{\min}[{\bf C}]=0$, which leads to $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$. ∎
Appendix J Proof of Proposition 3
@@
We have a lower bound of the objective function in (49) as ${\bf q}_{2}^{\top}{\bf B}{\bf q}_{2}\geq\nu_{\min}[{\bf B}]\|{\bf q}_{2}\|^{2}.$ From Conjecture 1, we have $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$ when $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$. Then, we have the lower bound as ${\bf q}_{2}^{\top}{\bf B}{\bf q}_{2}\geq(1-\alpha)\sigma_{2}^{2}\|{\bf q}_{2}\|^{2}.$ Here, ${\bf q}_{2}^{\star}=[0,...,0,\sqrt{\eta_{2}}]^{\top}$ satisfies the lower bound with $\|{\bf q}^{\star}_{2}\|^{2}=\eta_{2}$, which can be easily shown by the fact that all the entries in the last column and row of ${\bf B}$ are zeros except the last diagonal entry is $(1-\alpha)\sigma_{2}^{2}$ due to $f_{2,N}=0$ from Proposition 1. In other words, ${\bf q}_{2}^{\star}$ is an optimal solution of (49). We then have the optimal solution for (48) as ${\bf g}_{2}^{\star}={\bf Q}_{2,\text{sqrt}}{\bf q}_{2}^{\star}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$, since $f_{2,N}=0$ from Proposition 1. For the specific case of $N=3$, we have $\nu_{\min}[{\bf B}]=(1-\alpha)\sigma_{2}^{2}$ for any $\alpha$ (refer to Appendix I for the proof). Therefore, with the same analysis as above, ${\bf q}_{2}^{\star}=[0,0,\sqrt{\eta_{2}}]^{\top}$ (or equivalently ${\bf g}_{2}^{\star}={\bf Q}_{2,\text{sqrt}}{\bf q}_{2}^{\star}=[0,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$) is optimal in (49). ∎
Appendix K Proof of Corollary 1